i would like to use Matlab to solve this p.
Problem 2 Express in partial fraction form the signal 3s2 + 25 + 1 X(S) =- 93 +
Answer #1

*(S) = 33+ 251 3+5 35+25+1 s(5715 X®S) = + ce S 5+1 35725+) LA + Bste s(5+1) S 571 35+25+1 = ACS+D +s(83+C) compare (A+B) = 3XCO) = 5 + + mattab: functions pripik) - residue [[ 3213, [10103) are ii . P = Si [k] = [g Sa-i S=0 Emply implics at is propeoctave:3> [r,p,k]=residue([3 2 1],[1 0 1 01) 1 - li 1 + li 0 + li -0 - li 0 + Oi k = [] (0x0)